3 feet long 2 feet wide and 2 feet high what is the space of the dog house?
double the space?

Answer by richwmiller(17219)


3*2*2=12 ft^3
Double any one measurement and the space doubles.
3*2*4=24 ft^3
6*2*2=24 ft^3
3*4*2=24 ft^3